A basketball starts from rest and accelerates with an a㏄eleration of 0.395 m/s2 while moving down a 9.00 m long inclined plane, when it reaches the bottom, the ball rolls up another plane, where, after moving 15.00 m, it comes to rest. (a) What is the speed of the bahe bottom of the first plane (in m/s)? (Round your answer to at least two decimal places.) m/s (b) How long does it take to roll down the first plane (in s)?
Answer #1

(a) v^2 = vo^2 + 2 *a * d

= 0 + 2 * 0.395 * 9

v = 2.67 m/s (it is the speed of the ball at the bottom of the first plane )

(b) v = vo + a * t

2.67 = 0 + 0.395 t ---> t = 6.75 s